OpenStudy (shelby1290):

The manager of a hardware store sells batteries for $5 a package. She wants to see how much money she will earn if she increases the price in 10 cent increments.A model of the price change is the revenue function R(x)= -x^2+10x+3000, where x is the number of 10 cent increments and R(x) is in dollars. Determine the maximum revenue

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

@pooja195 @Nnesha

OpenStudy (anonymous):

@Englishguy

Directrix (directrix):

The graph of y= -x^2+10x+3000 is a parabola. I suspect that the vertex coordinates of the graph will give an idea of the number of ten cent increases that cause a maximum revenue.

OpenStudy (anonymous):

I'm no good with math, sorry

OpenStudy (anonymous):

i guess u get that from ur name

Directrix (directrix):

The maximum revenue should be the y-value of the vertex of the parabola.

Directrix (directrix):

To get the x coordinate of the vertex, crank this out: x vertex = -b/(2a) where b = 10 and a = -1. (These a and be values come from the coefficients of the x and x^2 terms of the revenue polynomial.)

Directrix (directrix):

@shelby1290 x vertex = -b/(2a) where b = 10 and a = -1. Question: